if 264g of N2H4 and 106g of N2O4 are mixed, calculate the number of moles of the excess reactant


1
Expert's answer
2021-10-30T01:15:03-0400

Molar mass of N2H4 = 32.0452 g/mol

264/32.0452= 8.24moles

Molar mass of N2O4= 92.011 g/mol

106/92.011=1.15moles

N2H4 + N2O4 → 3 N2 + 4 H2O

7(8.24)+7(1.15)

= 57.68+8.05=65.73moles